Calculate the amount of energy needed by a catapult to throw a stone of mass 500 g with a velocity of 10m/s


1
Expert's answer
2022-04-11T12:03:41-0400
"E=\frac{mv^2}{2}=\frac{0.5*10^2}{2}=25\:\rm J"